OpenStudy (anonymous):

child admission is $5.70 and adult admission is $9.20,175 tickets sold, total $1344.00 . How many adult tickets were sold that day?

OpenStudy (compassionate):

5.70c + 9.20a = 1,344.00 x + y = 175 Solve

OpenStudy (compassionate):

Or rather, c + a = 175 or clarity